• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Markowitz

    Log entry time 08:09:38 on November24,2008

    Entry number 250531

    Followups:

    keyword=Owl shift summary

    Started shift taking data. Beam trips were initially "high" (lots of grass on the bcm plots) but by 03:00 had become a bit better with ~3-5 minutes between trips.

    Run 3855 00:35 1M evts 0.03538C replay OK
    Run 3856 01:47 1M evts 0.03581C replay OK
    Run 3857 02:49 1M evts 0.0362C replay OK
    Run 3858 03:46 1M evts 0.03678C replay OK
    Run 3859 04:46 1M evts 0.03666C replay OK
    CODA froze 05:06, run 3860 had 100 DT, did not finish clean
    Run 3861 06:32 1M evts 0.03595C replay OK
    Run 3861 07:45 1M evts 0.03585C replay OK


    The small increasein the charge might indicate an increase in the deadtime per run. Did something change in the past day? Was a HV changed?